# Using the Jones Polynomial to Prove Infinite Families of Knots Satisfy the Cosmetic Surgery Conjecture

### v1

@@ 1 Introduction
For a knot $K$ to have cosmetic surgeries, the surgery slopes $r$ and $r^{\prime}$ must be $\pm 2$, the genus of $K$ must be $2$, and the Alexander polynomial of $K$ must be identically $1$ [DEL]. The knot $K$ must also be prime [Tao] and its Jones polynomial $V_{K}$ must satisfy $V_{K}^{\prime\prime}(1)=0$ and $V_{K}^{\prime\prime\prime}(1)=0$. [IchiharaWu]. The derivatives of Jones polynomials are examples of finite type invariants; the higher order derivatives are powerful but previous obstructions were more complicated relations dependent on surgery slopes [Ito]. We will derive, from [Ito] and the restrictions on the Alexander polynomial from [DEL], the following.
@@ 2 Relation to Previous Work
The knots considered here are knots of canonical genus $2$. The canonical genus of a knot is the minimal genus obtained from the Seifert algorithm over any possible diagram. While the potential difference between the genus and the canonical genus of any knot is unbounded [KobayashiKobayashi], considering knots of a given canonical genus is a useful way of ordering families of knots. Explicitly, all knots of canonical genus $2$ can be related by $\overline{t_{2}^{\prime}}$ moves to one of the knots: $5_{1}$, $6_{2}$, $6_{3}$, $7_{5}$, $_{6}, 77$$_{7}$, $8_{12}$, $8_{14}$, $8_{15}$, $9_{23}$, $9_{25}$, $9_{38}$, $9_{39}$, $9_{41}$, $10_{58}$, $_{97}, 1010$$_{101}$, $0_{120}, 11$$_{123}, 11$$_{148}, 11$$1_{329}$, $_{1097}, 1212$$_{1202}$, or $13_{4233}$ [Stoimenow]. A $\overline{t_{2}^{\prime}}$ move is the operation of changing a single crossing to three crossings, which corresponds to introducing a full twist to a band in a Seifert diagram, visualized as a collection of nodes and connecting bands. See 1.
@@
There is some overlap between canonical genus $2$ families of knots and knots previously shown to satisfy the Cosmetic Surgery Conjecture, but enumerating these overlaps verifies that the result is novel. For instance, the knots related to $5_{1}$ are precisely the genus $2$, 55-stranded pretzel knots. Pretzel knots have already been proven to satisfy the conjecture [StipsiczSzabo].
All prime alternating knots of genus $2$ and signature 0 (hence $\tau$ invariant 0) can be obtained by starting with one of six knots, $6_{3}$, $7_{7}$, $8_{12}$, $9_{41}$, $10_{58}$, $12_{1202}$, and performing $\overline{t_{2}^{\prime}}$ moves. (Proposition 3.2; [Stoimenow]) These are not excluded from consideration entirely, since it is possible to introduce crossing changes that do not preserve the alternating structure, but for part of the family the result has already been shown; [IchiharaJong] shows the result for a subset of these and [DEL] proves the result for alternating knots in general.
Previous work has proven many two bridge knots satisfy the conjecture [ichiharaSaito]. As all two bridge knots are alternating [Goodrick], there may be some overlap with each of the families of canonical genus $2$ knots, but these are limited as the particular collection of signs must begin as an alternating knot, and most of the choices of sign do not give alternating knots. Additionally, since the bridge number cannot decrease, when starting with a knot for which the bridge number is greater than $2$, the entire family will be disjoint.
The overlap with $3$ braid knots, which have been proven to satisfy the conjecture [Varvarezos], is finite, since if a $3$ braid knot has genus $2$ it must have $10$ or fewer crossings, so only a finite number of cases from each knot could have possible overlap.
Previous work [IchiharaJong] has shown some Montesinos knots satisfy the conjecture. According to the classifications of the paper, using their notation for Montesinos knots, all of their classes but (e2)-(e3) can be seen to be canonical genus $2$ knots, as shown below. However, there are many infinite families of canonical genus $2$ knots that are not Montesinos knots so this category is more general.
@@
The Montesinos knots not proven to lack cosmetic surgeries by [IchiharaJong] are a subset of the families generated by $6_{2}$, $7_{5}$, $_{6}, 77$$_{7}$ and $10_{58}$.
@@
(o1), $a,c,d\neq-1$, and $a,b,c,d\neq 0$. Proven by [IchiharaJong] to satisfy the conjecture for (o1’) where additionally $e$ is either 0 or 1-1.
@@
(o3), $b,c\neq-1,0$, $a\neq 0$. Proven by [IchiharaJong] to satisfy the conjecture for (o3’), where additionally $|a|=1$.
@@
(o4) $b,c,d\neq-1,0$, $a\neq 0$. Proven by [IchiharaJong] to satisfy the conjecture for (o4’), where $|a|=1$.
@@
is not of canonical genus $2$ and has been proven to satisfy the conjecture [IchiharaJong]. See figure 8. Comparison with every knot family of canonical genus $2$ does not give any matches. Regardless, if there was overlap, counting parameters shows that (e2) would be a proper subset of any knot family of canonical genus $2$, as the smallest number of parameters that any family of canonical genus $2$ has is $4$.
@@
is not of canonical genus $2$ and has been proven to satisfy the conjecture [IchiharaJong]. See figure 9. Comparison gives no matches, but again, if there was overlap, it would be at most a proper subset of any canonical genus $2$ family, as it only has one parameter.
@@
A final restriction can be obtained by evaluating the Jones polynomial itself at a fifth root of unity, $\xi_{5}$. Using the results of [Detcherry], if a knot $K$ has a purely cosmetic surgery the slope of the surgery will be of the form $\frac{1}{5k}$ unless $J_{k}(\xi_{5})=1$. Hence as $2$ is not of this form, if $J_{K}(\xi_{5})\neq 1$, we may exclude the possibility of cosmetic surgery. In this context, this will give a mod $5$ congruence restriction on the number of pairs of twists in each possible location. Although this will not be needed in this computation, it works very well with these methods and is a potentially useful way of identifying infinitely many knots within a family which satisfy the conjecture.
These results build upon other foundational work, including [GL] [OS] [Wang] [Wu] [NiWu] [Hanselman] [FPS]. Additionally, Ren has proven [Ren] that only hyperbolic knots could have cosmetic surgeries, excluding the case of satellite knots.
@@ 3 Jones Polynomial Restrictions
Now, considering the result of [Ito] we can consider the finite type invariant of one degree higher, which gives a relation between the fourth derivative of the Jones polynomial evaluated at $1$ and the Alexander polynomial, and use the results of [DEL] to prove the position stated in the introductory section, that if $K$ has cosmetic surgeries, $V_{K}^{\prime\prime\prime\prime}(1)=0$.
@@
From Corollary 1.5(i) of [Ito], we have
@@
Additionally, by Lemma 2.1 of [Ito] we have
@@
where $a_{2i}(K)$ is the coefficient of $z^{2i}$ in the Conway polynomial and $j_{n}$ is the coefficient of $h^{n}$ in the Jones polynomial $V_{K}(e^{h})$. However, if $K$ has a cosmetic surgery, the Alexander polynomial and hence the Conway polynomial must be trivial [DEL]. Therefore, we can simplify these formulae to:

### v2

@@ 1 Introduction
For a knot $K$ to have cosmetic surgeries, the surgery slopes $r$ and $r^{\prime}$ must be $\pm 2$, the genus of $K$ must be $2$, and the Alexander polynomial of $K$ must be identically $1$ [1]. The knot $K$ must also be prime [17] and its Jones polynomial $V_{K}$ must satisfy $V_{K}^{\prime\prime}(1)=0$ and $V_{K}^{\prime\prime\prime}(1)=0$. [9]. The derivatives of Jones polynomials are examples of finite type invariants; the higher order derivatives are powerful but previous obstructions were more complicated relations dependent on surgery slopes [10]. We will derive, from [10] and the restrictions on the Alexander polynomial from [1], the following.
@@ 2 Relation to Previous Work
The knots considered here are knots of canonical genus $2$. The canonical genus of a knot is the minimal genus obtained from the Seifert algorithm over any possible diagram. While the potential difference between the genus and the canonical genus of any knot is unbounded [11], considering knots of a given canonical genus is a useful way of ordering families of knots. Explicitly, all knots of canonical genus $2$ can be related by $\overline{t_{2}^{\prime}}$ moves to one of the knots: $5_{1}$, $6_{2}$, $6_{3}$, $7_{5}$, $_{6}, 77$$_{7}$, $8_{12}$, $8_{14}$, $8_{15}$, $9_{23}$, $9_{25}$, $9_{38}$, $9_{39}$, $9_{41}$, $10_{58}$, $_{97}, 1010$$_{101}$, $0_{120}, 11$$_{123}, 11$$_{148}, 11$$1_{329}$, $_{1097}, 1212$$_{1202}$, or $13_{4233}$ [16]. A $\overline{t_{2}^{\prime}}$ move is the operation of changing a single crossing to three crossings, which corresponds to introducing a full twist to a band in a Seifert diagram, visualized as a collection of nodes and connecting bands. See 1.
@@
There is some overlap between canonical genus $2$ families of knots and knots previously shown to satisfy the Cosmetic Surgery Conjecture, but enumerating these overlaps verifies that the result is novel. For instance, the knots related to $5_{1}$ are precisely the genus $2$, 55-stranded pretzel knots. Pretzel knots have already been proven to satisfy the conjecture [15].
All prime alternating knots of genus $2$ and signature 0 (hence $\tau$ invariant 0) can be obtained by starting with one of six knots, $6_{3}$, $7_{7}$, $8_{12}$, $9_{41}$, $10_{58}$, $12_{1202}$, and performing $\overline{t_{2}^{\prime}}$ moves. (Proposition 3.2; [16]) These are not excluded from consideration entirely, since it is possible to introduce crossing changes that do not preserve the alternating structure, but for part of the family the result has already been shown; [7] shows the result for a subset of these and [1] proves the result for alternating knots in general.
Previous work has proven many two bridge knots satisfy the conjecture [8]. As all two bridge knots are alternating [4], there may be some overlap with each of the families of canonical genus $2$ knots, but these are limited as the particular collection of signs must begin as an alternating knot, and most of the choices of sign do not give alternating knots. Additionally, since the bridge number cannot decrease, when starting with a knot for which the bridge number is greater than $2$, the entire family will be disjoint.
The overlap with $3$ braid knots, which have been proven to satisfy the conjecture [18], is finite, since if a $3$ braid knot has genus $2$ it must have $10$ or fewer crossings, so only a finite number of cases from each knot could have possible overlap.
Previous work [7] has shown some Montesinos knots satisfy the conjecture. According to the classifications of the paper, using their notation for Montesinos knots, all of their classes but (e2)-(e3) can be seen to be canonical genus $2$ knots, as shown below. However, there are many infinite families of canonical genus $2$ knots that are not Montesinos knots so this category is more general.
@@
The Montesinos knots not proven to lack cosmetic surgeries by [7] are a subset of the families generated by $6_{2}$, $7_{5}$, $_{6}, 77$$_{7}$ and $10_{58}$.
@@
(o1), $a,c,d\neq-1$, and $a,b,c,d\neq 0$. Proven by [7] to satisfy the conjecture for (o1’) where additionally $e$ is either 0 or 1-1.
@@
(o3), $b,c\neq-1,0$, $a\neq 0$. Proven by [7] to satisfy the conjecture for (o3’), where additionally $|a|=1$.
@@
(o4) $b,c,d\neq-1,0$, $a\neq 0$. Proven by [7] to satisfy the conjecture for (o4’), where $|a|=1$.
@@
is not of canonical genus $2$ and has been proven to satisfy the conjecture [7]. See figure 8. Comparison with every knot family of canonical genus $2$ does not give any matches. Regardless, if there was overlap, counting parameters shows that (e2) would be a proper subset of any knot family of canonical genus $2$, as the smallest number of parameters that any family of canonical genus $2$ has is $4$.
@@
is not of canonical genus $2$ and has been proven to satisfy the conjecture [7]. See figure 9. Comparison gives no matches, but again, if there was overlap, it would be at most a proper subset of any canonical genus $2$ family, as it only has one parameter.
@@
A final restriction can be obtained by evaluating the Jones polynomial itself at a fifth root of unity, $\xi_{5}$. Using the results of [2], if a knot $K$ has a purely cosmetic surgery the slope of the surgery will be of the form $\frac{1}{5k}$ unless $J_{k}(\xi_{5})=1$. Hence as $2$ is not of this form, if $J_{K}(\xi_{5})\neq 1$, we may exclude the possibility of cosmetic surgery. In this context, this will give a mod $5$ congruence restriction on the number of pairs of twists in each possible location. Although this will not be needed in this computation, it works very well with these methods and is a potentially useful way of identifying infinitely many knots within a family which satisfy the conjecture.
These results build upon other foundational work, including [5] [13] [19] [20] [12] [6] [3]. Additionally, Ren has proven [14] that if the Cosmetic Surgery Conjecture is false, there must exist a hyperbolic knot which has cosmetic surgeries.
@@ 3 Jones Polynomial Restrictions
Now, considering the result of [10] we can consider the finite type invariant of one degree higher, which gives a relation between the fourth derivative of the Jones polynomial evaluated at $1$ and the Alexander polynomial, and use the results of [1] to prove the position stated in the introductory section, that if $K$ has cosmetic surgeries, $V_{K}^{\prime\prime\prime\prime}(1)=0$.
@@
From Corollary 1.5(i) of [10], we have
@@
Additionally, by Lemma 2.1 of [10] we have
@@
where $a_{2i}(K)$ is the coefficient of $z^{2i}$ in the Conway polynomial and $j_{n}$ is the coefficient of $h^{n}$ in the Jones polynomial $V_{K}(e^{h})$. However, if $K$ has a cosmetic surgery, the Alexander polynomial and hence the Conway polynomial must be trivial [1]. Therefore, we can simplify these formulae to: